Markov Chains

Consider the Markov chain with transition matrix P = [ 0 1

1 0].

1) Compute several powers of P by hand. What do you notice?

2) Argue that a Markov chain with P as its transition matrix cannot stabilize unless both initial probabilities are 1/2.
Answer #1

0 1 1 0 P =

1) Observe that P^2 =I, hence the order of P is 2. Even powers of P are I and odd powers of P is P.

2) The Markov chain stabilizes when T

We need to find T P2 such that \pi (P-I) = 0 and p_1+p_2=1

-1 1 T 1 0 -1

Hence the solution is P P. But p_1+p_2=1, therefore P1 P2 = 1/2 is the only solution.